PFO operational status:
IDA engineers visited the station and replaced the PFO-00 Q330 digitizer.
IDA engineers visited the station and replaced the STS-1s with an STS-6 seismometer, and moved the episensor to the wellhead by the new PFO-00 Q330. A new gateway was added for a second link between UCSD and the station. Station location metadata has been updated to reflect the location of the STS-6 seismometer.
An IDA engineer vistited the station and replaced the PFO-10 Q330 digitizer.
IDA engineers visited the station and replaced the Trillium 240 seismometer at 2017,131-16:15.
An IDA engineer visited the station and replaced the STS-1V and STS-1N seismometers.
The background noise on channels from the STS-1 was eliminated after IDA engineers removed an unnecessary ground between power and vault equipment on day 2014,069.
All channels from the STS-1 are showing high background noise with 30s pulses since day 2013,206 (July25).
An IDA engineer visited the station and replaced the STS-1 North-South seismometer.
IDA engineers visited the station on day 2012,225 to investigate the noise problem for the STS-1 N-S sensor. Despite changing cables to the sensor, the base plate, and reinstalling the sensor the noise problem continues. Plans to replace the STS-1 are underway.
An IDA engineer visited the station to investigate the noise problem with the STS-1 N-S seismometer, and to install a bi-directional GPS cable between digitizers.
IDA engineers visited the station and upgraded the power system and replaced the North-South STS-1, and GPS antenna.
An IDA engineer visited the station and replaced the base plate for the STS-1V, and re-installed the seismometer. The data appears to have improved considerably, but will continue to be evaluated for future action.
An IDA engineer visited the station and replaced the E300 and re-evacuated the bell jar for the STS-1. The long period pulses on the vertical channels from the STS-1 disappeared for a while but later reappeared. Troubleshooting is still in progress.
There are long period pulses on the vertical channels from the STS-1 seismometer since day 2010,286, for reasons yet to be determined.
An IDA engineer visited the station and replaced the STS-1N seismometer on day 2010,211 (July 30).
The North-South channels from the STS-1 seismometer are showing high background noise since the seismometer interface box and cables were changed on day 2010,132 (May 12). IDA engineers are investigating the problem.
The vertical and horizontal channels from the Episensor accelerometer were non-seismic from about day 2008,142 to 2008,339 due to a faulty sensor. The channels were restored after an IDA engineer visited the station and replaced the sensor.
IDA station PFO has been upgraded to a Quanterra Q330HR system.
An IDA engineer visited the station and replaced the STS-2 sensor on day 2007,324 (Nov. 20, 2007).
Data for the episensor strong motion detector are seismic again as of about day 342-23:30 after an IDA engineer reconfigured the cabling to the episensor.
An IDA engineer visited the site and replaced the Power Supply in the recording room.
The broadband channels from the STS-1 sensor had excessive noise above 1 Hz from about day 270 – 285 (9/27 to 10/12) period due to faulty grounding of station equipment. The problem was resolved after an IDA engineer visited the station and reconfigured the Main power supply.
The short period channels from the STS-2 sensor had excessive background noise from day 194 to 244 (July 13 – Sept.1).
Data quality for all channels from the episensor strong motion detector is poor for reasons yet to be determined.
IDA station PFO has been upgraded and moved to a new vault. The MK-8 system includes a new Episensor strong motion detector and microbarometer in addition to STS-1 and STS-2 sensors. Data acquisition for PFO switched from the old digitizer to the new one at about 194-00:00.
The microbarometer was replaced by an IDA engineer and the wdi channel was restored.
The wdi channel is out due to a faulty microbarometer.
Telemetry to the station was down day 289-19:13 (Oct.16) to day 303. When telemetry was restored normal data acquisition resumed.
The East-West channels from the STS-1 sensor are normal as of day 352 (12/17/04) after the STS-1 sensor’s masses were recentered. The station was down from day 12/30/04 to 01/02/05 after the station lost power during a storm. The station came back up after an IDA engineer power cycled the DAS.
The East-West channels from the STS-1 sensor are saturated because the sensor’s masses need to be recentered (12/10/2004).
Station is up as of (8/20/2004) after IDA engineers replaced ethernet switches in the vault and recording trailer, which were damaged by the lightning strike. There is no data for the time period from day 226-19:00 to 233-17:12.
Station experienced lightning strike on 8/13/2004. Telemetry down and station visit scheduled to assess station status.
The mass recentering seems to have addressed the STS2 data problem.
The amplitudes of STS2 data start to look anomalous – perhaps due to mass position problem.
Station operational again after an IDA engineer replaced the ethernet switch.
Station down due to a bad ethernet switch. Network connection to the station’s digitizer was lost as of 2003:235-23:47 (August 23,2003).
New main power supply installed.
New station processor installed.
An IDA engineer visited the station today to calibrate the FBA and reset the free periods of the STS1 sensors.
The station is operating normally. Today the data stream from the vertical channel of FBA was repaired.